Jane is walking at such speed o that her kinetic energy is 20 J. What is her speed? It started to rain. Jane runs at a rate of 4.2 m/s. Find the change in her kinetic energy. Jane has a mass of 50 kg.


1
Expert's answer
2022-01-17T09:58:55-0500

The kinetic energy is given as follows:


"K_1 = \dfrac{mv_1^2}{2}"

Thus, the speed is:


"v_1 = \sqrt{\dfrac{2K_1}{m}}=\sqrt{\dfrac{2\cdot 20J}{50kg}} \approx 0.89m/s"

In the second case:


"K_2 = \dfrac{mv^2_2}{2} = \dfrac{50\cdot 4.2^2}{2} = 441J"

Thus, the change in kinetic energy is:


"K_2-K_1 = 441J-20J = 421J"

Answer. 0.89m/s, 421J.